40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12083     Entry time: Tue Apr 19 18:37:29 2016     In reply to: 12079     Reply to this: 12085   12089
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: Laser swap + optical layout 

Summary of work done over the last two days

  1. Lightwave NPRO + controller moved to PSL table
    • ​​The interlock is not connected to the controller
    • Controller is not powered
  2. Innolight NPRO + controller installed at endtable
    • ​​​​Interlock has been connected
    • For initial alignment purposes, I'm running it at an injection current of 1.000A (~50mW of IR out of the NPRO)
    • Temperature of crystal set to 31.66 degrees in anticipation of operation in the nominal state
  3. Laying out optics
    • ​​Given that the mode out of the NPRO is different from that from the Lightwave, the mode-matching had to be re-done
    • Attachment #1 shows the mode-matching solution being implemented
    • Current state - I've placed all the optics up to and including the doubling crystal + oven. Alignment through IR Faraday is pretty good, QWP+HWP angles optimized to maximize transmission through the Faraday (<10% loss). Oven has been hooked up to temperature controller, and is currently set to 36.3 degrees. Coarse alignment into doubling crystal done at lower power. Even with the low IR power, I am able to see some green. It remains to turn the injection current up and do the fine alignment + lens position tweaking to maximize the green power from the doubling crystal - with ~1W of power, assuming 2%/W SHG efficiency, we should be seeing 20 mW of green (which is probably way too much)

Immediate next steps:

  1. Some optimization to be done with regards to beam dumps for rejected beam from IR Faraday. Also double check to make sure that the reflected beam from L1 doesn't go back directly to the laser (at the moment it doesn't, is there a standard way to do this? I was trying to have the lens as close to normal incidence as possible, but I may not have been entirely successful which is why the reflected beam is not going straight back at the moment).
  2. Optimize mode-matching into the doubling crystal
  3. Once the desired green mode is obtained, continue with the rest of the layout
  4. Update CAD drawing to reflect new layout

 

Attachment 1: IR_modematch_19April2016_2.pdf  106 kB  | Hide | Hide all
IR_modematch_19April2016_2.pdf
ELOG V3.1.3-